Solve:

"Determin "k" if -2 is a root of:
2x^2+ky+2=0

Answer by stanbon(75887)

"Determine "k" if -2 is a root of:
2x^2+kx+2=0
-------------
If -2 is a root, f(-2)=0
f(x)=2x^2+kx+2
f(-2)=2(-2)^2+k(-2)+2=0
8-2k+2=0
-2k=-10
k=5
============
Cheers,
Stan H.